# Classical harmonic analysis viewed through the prism of noncommutative geometry

### v4

@@ 8 Examples of triples and Banach Fredholm modules
see, e.g., [Haa06, Proposition 8.4.1 p. 228].
@@
Let $1<p<\infty$ and let $D\overset{\mathrm{def}}{=}\frac{1}{\mathrm{i}}\frac{\mathrm{d}}{\mathrm{d}x}$ with domain $\operatorname{dom}D=\mathrm{W}^{1,p}(\mathbb{R})$. Then
@@
Proof : If $f\in\mathrm{W}^{1,\infty}(\mathbb{R})$ then $M_{f}(\mathrm{W}^{1,p}(\mathbb{R}))\subset\mathrm{W}^{1,p}(\mathbb{R})$ by [BeH21, Theorem 4.1 p. 284]. Moreover, if $u\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$ we have
$D,M_{f}]u=(DM_{f}-M_{f})u=\tfrac{1}{\mathrm{i}}\big((fu)^{\prime}-fu^{\prime}\big)=\tfrac{1}{\mathrm{i}}f^{\prime}u.$
Hence $D,M_{f}]=\tfrac{1}{\mathrm{i}}M_{f^{\prime}}$ extends boundedly on $\mathrm{L}^{p}(\mathbb{R})$ with $\left\|[D,M_{f}]\right\|_{\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})}=\left\|f^{\prime}\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}$.
Let $(T_{t})_{t\in\mathbb{R}}$ be the translation group on $\mathrm{L}^{p}(\mathbb{R})$ given by $(T_{t}u)(x)=u(x+t)$. Then for any $t\in\mathbb{R}$ we have $T_{t}=\mathrm{e}^{\mathrm{i}tD}$ (see [EnN00, Proposition 1 p. 66] or [Haa06, Proposition 8.4.2 p. 229]) and
(8.22) $T_{t}M_{f}T_{-t}=M_{f(\cdot+t)},\qquad t\in\mathbb{R}.$
Let $f\in\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{R}))$. Since $D,M_{f}]\in\mathrm{B}(\mathrm{L}^{p}(\mathbb{R}))$ by Proposition 8.2 we deduce that
[…]
If $|t|$ is small, we infer that
[…]
$\leqslant\int_{0}^{1}\left\|T_{st}(\mathrm{i}[D,M_{f}])T_{-st}\right\|_{\mathrm{B}(\mathrm{L}^{p}(\mathbb{R}))}\mathop{}\mathopen{}\mathrm{d}s\leqslant\left\|[D,M_{f}]\right\|_{\mathrm{B}(\mathrm{L}^{p}(\mathbb{R}))}.$
By Banach–Alaoglu, there exist $t_{\alpha}\to 0$ and $g\in\mathrm{L}^{\infty}(\mathbb{R})$ such that $\Delta_{t_{\alpha}}f\overset{}{\longrightarrow}g$ in the weak* topology of the dual space $\mathrm{L}^{\infty}(\mathbb{R})$, with $\left\|g\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}\leqslant\left\|[D,M_{f}]\right\|_{\mathrm{B}(\mathrm{L}^{p}(\mathbb{R}))}$. For any function $\varphi\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$ we have
$\int_{\mathbb{R}}(\Delta_{t_{\alpha}}f)\varphi=\int_{\mathbb{R}}f\frac{\varphi(\cdot-t_{\alpha})-\varphi}{t_{\alpha}}\longrightarrow-\int_{\mathbb{R}}f\,\varphi^{\prime}$
since $\frac{\varphi(\cdot-t)-\varphi}{t}\to-\varphi^{\prime}$ in $\mathrm{L}^{1}(\mathbb{R})$ as $t\to 0$. Passing to the weak* limit gives
$\int_{\mathbb{R}}g\varphi=-\int_{\mathbb{R}}f\varphi^{\prime},\quad\varphi\in\mathrm{C}_{c}^{\infty}(\mathbb{R}),$
i.e., $g$ is the weak derivative of $f$. Hence $f\in\mathrm{W}^{1,\infty}(\mathbb{R})$ and $f^{\prime}=g\in\mathrm{L}^{\infty}(\mathbb{R})$.
@@
Proof : First proof. Consider the infinitesimal generator $A\overset{\mathrm{def}}{=}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ of the left translation group $(T_{t})_{t\in\mathbb{R}}$ on the Banach space $\mathrm{L}^{p}(\mathbb{R})$. By the Laplace formula [HvNVW18, Proposition G.4.1 p. 532] for the resolvent, we have for $\operatorname{Re}\lambda>0$ the equality
@@
Second proof. The resolvent operator $(\mathrm{i}+D)^{-1}\colon\mathrm{L}^{p}(\mathbb{R})\to\mathrm{W}^{1,p}(\mathbb{R})$ is bounded by definition. Let $f\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$. Consider a bounded open subset $\Omega\subset\mathbb{R}$ such that $\operatorname{supp}f\subset\Omega$. Note that we have a bounded operator $M_{f}\colon\mathrm{W}^{1,p}(\mathbb{R})\to\mathrm{W}^{1,p}(\Omega)$. We denote by $J\colon\mathrm{W}^{1,p}(\Omega)\to\mathrm{L}^{p}(\Omega)$ the canonical injection, which is compact by [Eva10, Remark p. 289]. If $\mathcal{E}\colon\mathrm{L}^{p}(\Omega)\to\mathrm{L}^{p}(\mathbb{R})$ is the extension map by 0, then the operator $M_{f}(\mathrm{i}+D)^{-1}$ admits the factorization
$M_{f}(\mathrm{i}+D)^{-1}=E\circ J\circ M_{f}\circ(\mathrm{i}+D)^{-1},$
Actually, we have the following quantitative compactness result relying on the Poincaré inequality [EdE23, Theorem 2.4 p. 21]
for a bounded convex subset $\Omega$ of $\mathbb{R}^{d}$.
Proposition 8.15
Suppose that $1<p<\infty$. For any function $f\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$ the operator $M_{f}(D+\mathrm{i})^{-1}$ belongs to $S_{\mathrm{app}}^{1,\infty}(\mathrm{L}^{p}(\mathbb{R}))$.
Proof : Since $D^{2}=-\Delta$, we have the algebraic identity
$(D+\mathrm{i})^{-1}=-(\mathrm{i}-D)(1-\Delta)^{-1}.$
Note that the Bessel potential operator $(1-\Delta)^{-1}$ induces a bounded operator $(1-\Delta)^{-1}\colon\mathrm{L}^{p}(\mathbb{R})\to\mathrm{W}^{2,p}(\mathbb{R})$, while the operator $\mathrm{i}-D\colon\mathrm{W}^{2,p}(\mathbb{R})\to\mathrm{W}^{1,p}(\mathbb{R})$ is bounded by [Kry08, Exercise 8 p. 10]. Hence the operator
$S\overset{\mathrm{def}}{=}(\mathrm{i}-D)(1+\Delta)^{-1}\colon\mathrm{L}^{p}(\mathbb{R})\to\mathrm{W}^{1,p}(\mathbb{R})$
is bounded. Let $I$ be a fixed bounded interval of $\mathbb{R}$ containing $\operatorname{supp}f$. We consider the restriction operator $R_{I}\colon\mathrm{W}^{1,p}(\mathbb{R})\to\mathrm{W}^{1,p}(I)$ and the extension by zero operator $E_{I}\colon\mathrm{L}^{p}(I)\to\mathrm{L}^{p}(\mathbb{R})$. If $J\colon\mathrm{W}^{1,p}(I)\to\mathrm{L}^{p}(I)$ is the Sobolev embedding on the bounded interval $I$, then we have the factorization
(8.28) $M_{f}(D+\mathrm{i})^{-1}=E_{I}\underbrace{M_{f}|_{\mathrm{L}^{p}(I)}}_{M_{f}^{I}\overset{\mathrm{def}}{=}}\underbrace{J}_{\mathrm{W}^{1,p}(I)\hookrightarrow\mathrm{L}^{p}(I)}\underbrace{R_{I}S}_{S_{I}\overset{\mathrm{def}}{=}}.$
Consider a partition of $I$ into $n$ subintervals $I_{1},\ldots,I_{n}$ of equal length $\frac{|I|}{n}$. Consider the linear operator $P_{n}\colon\mathrm{W}^{1,p}(I)\to\mathrm{L}^{p}(I)$ defined by
$(P_{n}g)(x)\overset{\mathrm{def}}{=}\sum_{k=1}^{n}\bigg(\frac{1}{|I_{k}|}\int_{I_{k}}g(t)\mathop{}\mathopen{}\mathrm{d}t\bigg)1_{I_{k}}(x),\qquad g\in\mathrm{W}^{1,p}(I).$
Observe that $\operatorname{rank}P_{n}\leqslant n$. On each subinterval $I_{k}$, the Poincaré inequality yields
[…]
Summing over $k=1,\dots,n$ (the $I_{k}$ are disjoint) gives
$\left\|g-P_{n}(g)\right\|_{\mathrm{L}^{p}(I)}^{p}=\sum_{k=1}^{n}\left\|g-\frac{1}{|I_{k}|}\int_{I_{k}}g(t)\mathop{}\mathopen{}\mathrm{d}t\right\|_{\mathrm{L}^{p}(I_{k})}^{p}\leqslant\bigg(\frac{|I|}{n}\bigg)^{p}\sum_{k=1}^{n}\left\|g^{\prime}\right\|_{\mathrm{L}^{p}(I_{k})}^{p}=\bigg(\frac{|I|}{n}\bigg)^{p}\left\|g^{\prime}\right\|_{\mathrm{L}^{p}(I)}^{p}.$
Taking the $p$-th root, we obtain
$\left\|g-P_{n}g\right\|_{\mathrm{L}^{p}(I)}\leqslant C\frac{|I|}{n}\left\|g^{\prime}\right\|_{\mathrm{L}^{p}(I)}\leqslant C\frac{|I|}{n}\left\|g\right\|_{\mathrm{W}^{1,p}(I)}.$
Hence, we obtain the estimate
$\left\|J-P_{n}\right\|_{\mathrm{W}^{1,p}(I)\to\mathrm{L}^{p}(I)}\leqslant C\frac{|I|}{n}.$
By the ideal property of approximation numbers, this implies
[…]
From the previous factorization and the ideal property of approximation numbers,
[…]
We conclude with (5.9).
@@
and can be seen by [HvNVW16, Proposition 5.2.2 p. 389] as a Fourier multiplier with symbol $\mathrm{i}\,\operatorname{\mathrm{sign}}(\xi)$, i.e., we have
(8.32) $\widehat{\mathcal{H}(f)}(\xi)=-\mathrm{i}\,\operatorname{\mathrm{sign}}(\xi)\hat{f}(\xi),\quad f\in\mathcal{S}(\mathbb{R}),\xi\in\mathbb{R}.$
@@
In this section, we will use the classical operators
$\partial_{\overline{z}}\overset{\mathrm{def}}{=}\frac{1}{2}\bigg(\frac{\partial}{\partial x}+\mathrm{i}\frac{\partial}{\partial y}\bigg)\quad\text{and}\quad\partial_{z}\overset{\mathrm{def}}{=}\frac{1}{2}\bigg(\frac{\partial}{\partial x}-\mathrm{i}\frac{\partial}{\partial y}\bigg).$
Recall that $\Delta=4\partial_{\overline{z}}\partial_{z}$ The operators $\partial_{\overline{z}}$, $\partial_{z}$ and $\Delta$ can be seen as Fourier multipliers on with symbol $\mathrm{i}\pi\zeta$, $\mathrm{i}\pi\overline{\zeta}$ and $|\zeta|^{2}$, see [AIM09, p. 99]. In the sequel, we consider the Dirac operator
(8.39) $D\overset{\mathrm{def}}{=}\begin{bmatrix}0&-2\partial_{\overline{z}}\\ 2\partial_{z}&0\\ \end{bmatrix}$.
acting on the subspace $\mathrm{W}^{1,p}(\mathbb{C})\oplus\mathrm{W}^{1,p}(\mathbb{C})$ of the Banach space $\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})$. The square of this operator is given by $D^{2}=\begin{bmatrix}-4\partial_{\overline{z}}\partial_{z}&0\\ 0&-4\partial_{z}\partial_{\overline{z}}\\ \end{bmatrix}=\begin{bmatrix}-\Delta&0\\ 0&-\Delta\\ \end{bmatrix}$. In the sequel, we use the notations $\partial\overset{\mathrm{def}}{=}2\partial_{z}$ and $\partial^{*}\overset{\mathrm{def}}{=}-2\partial_{\overline{z}}$. See [LaM89, p. 119] for more information on this operator. We can introduce the isometric homomorphism $\pi\colon\mathrm{L}^{\infty}(\mathbb{C})\to\mathrm{B}(\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C}))$, $f\mapsto\begin{bmatrix}M_{f}&0\\ 0&M_{f}\\ \end{bmatrix}$.
We begin by determining the Lipschitz algebra.
@@
Suppose that $1<p<\infty$. Consider the operator $D\overset{\mathrm{def}}{=}\begin{bmatrix}0&-2\partial_{\overline{z}}\\ 2\partial_{z}&0\end{bmatrix}$ with domain $\operatorname{dom}D=\mathrm{W}^{1,p}(\mathbb{C})\oplus\mathrm{W}^{1,p}(\mathbb{C})$. Then
@@
(8.41) $D,f]=\begin{bmatrix}0&-2M_{\partial_{\overline{z}}f}\\ 2M_{\partial_{z}f}&0\end{bmatrix}$
@@
Proof : If $f\in\mathrm{W}^{1,\infty}(\mathbb{C})$, then by then $M_{f}(\mathrm{W}^{1,p}(\mathbb{C}))\subset\mathrm{W}^{1,p}(\mathbb{C})$. Hence $\operatorname{dom}D$ is invariant under the operator $\pi(f)$. For any $u,v\in\mathrm{C}_{c}^{\infty}(\mathbb{C})$, we have
(8.43) $\big[D,f\big]=\begin{bmatrix}0&-2\partial_{\overline{z}}\\ 2\partial_{z}&0\\ \end{bmatrix}\begin{bmatrix}M_{f}&0\\ 0&M_{f}\\ \end{bmatrix}-\begin{bmatrix}M_{f}&0\\ 0&M_{f}\\ \end{bmatrix}\begin{bmatrix}0&-2\partial_{\overline{z}}\\ 2\partial_{z}&0\\ \end{bmatrix}$
$=\begin{bmatrix}0&-2\partial_{\overline{z}}M_{f}\\ 2\partial_{z}M_{f}&0\\ \end{bmatrix}+\begin{bmatrix}0&2M_{f}\partial_{\overline{z}}\\ -2M_{f}\partial_{z}&0\\ \end{bmatrix}=\begin{bmatrix}0&-2[\partial_{\overline{z}},M_{f}]\\ 2[\partial_{z},M_{f}]&0\end{bmatrix}$.
Consequently, we have
@@
Hence, we obtain (8.41). Since the functions $\partial_{\overline{z}}f$ and $\partial_{z}f$ belong to $\mathrm{L}^{\infty}(\mathbb{C})$, we infer that this operator admits a bounded extension. We deduce that $\mathrm{W}^{1,\infty}(\mathbb{C})\subset\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{C}))$. The inequality $\big\|[D,f]\big\|\leqslant 2\max\Big\{\left\|\partial_{z}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})},\left\|\partial_{\overline{z}}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})}\Big\}$ follows from a simple computation:
@@
Now, suppose that $f\in\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{C}))$. By definition, the operator $\pi(f)$ leaves $\operatorname{dom}D=\mathrm{W}^{1,p}(\mathbb{C})\oplus\mathrm{W}^{1,p}(\mathbb{C})$ invariant and $D$ induces a bounded operator on $\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})$. Writing
$D,f]=\begin{bmatrix}0&-2[\partial_{\overline{z}},M_{f}]\\ 2[\partial_{z},M_{f}]&0\end{bmatrix}$,
we see that $\partial_{z},M_{f}]$ and $\partial_{\overline{z}},M_{f}]$ are bounded on $\mathrm{L}^{p}(\mathbb{C})$. Since $\partial_{x}=\partial_{z}+\partial_{\overline{z}}$ and $\partial_{y}=\tfrac{1}{\mathrm{i}}\,(\partial_{z}-\partial_{\overline{z}})$, it follows that $\partial_{x},M_{f}]$ and $\partial_{y},M_{f}]$ are bounded on $\mathrm{L}^{p}(\mathbb{C})$. Fix $j\in\{1,2\}$ and apply Proposition 8.2 with $R=M_{f}$ and the one-parameter translation group $(\tau_{te_{j}})_{t\in\mathbb{R}}$ in direction $e_{j}$ with generator $\partial_{j}$. We obtain, for all small $t\neq 0$,
@@
Consequently, we have
$\big\|\Delta_{t}^{(j)}f\big\|_{\mathrm{L}^{\infty}(\mathbb{C})}=\big\|M_{\Delta_{t}^{(j)}f}\big\|_{\mathrm{L}^{p}(\mathbb{C})\to\mathrm{L}^{p}(\mathbb{C})}=\left\|\int_{0}^{1}\tau_{ste_{j}}[\partial_{x_{j}},M_{f}]\,\tau_{-ste_{j}}\mathop{}\mathopen{}\mathrm{d}s\right\|_{\mathrm{L}^{p}(\mathbb{C})\to\mathrm{L}^{p}(\mathbb{C})}$

### v5

@@ 8.2 The Dirac operator on $\ell^{p}_{N}(\mathrm{L}^{p}(\mathbb{R}^{d}))$
Assume that $d=1$. We have $N=2^{\lfloor\frac{d}{2}\rfloor}=1$. Then the Dirac operator is given by
$D=-\mathrm{i}\gamma_{1}\otimes\partial_{x}=\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x},$
with domain $\mathrm{W}^{1,p}(\mathbb{R})$ acting on $\mathrm{L}^{p}(\mathbb{R}))$. Here $\gamma_{1}=[1]$ is an $N\times N$ selfadjoint matrix satisfying the Clifford relation $\gamma_{1}^{2}=\mathrm{Id}_{\mathbb{C}^{N}}$. In the scalar one-dimensional case we take $N=1$ and choose the representation of the Clifford algebra given by
$\gamma_{1}=1\in\mathrm{M}_{1}(\mathrm{C}).$
@@
see, e.g., [Haa06, Proposition 8.4.1 p. 228]. The following result is a consequence of Proposition 8.13.
@@
Suppose that $1<p<\infty$. Let $D\overset{\mathrm{def}}{=}\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ with domain $\operatorname{dom}D=\mathrm{W}^{1,p}(\mathbb{R})$. Then
@@
The following result is a particular case of Proposition 8.14. We give an alternative proof.
@@
The next result is a particular case of Proposition 8.21. We give another argument.
@@
Proof : Consider the infinitesimal generator $A\overset{\mathrm{def}}{=}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ of the left translation group $(T_{t})_{t\in\mathbb{R}}$ on the Banach space $\mathrm{L}^{p}(\mathbb{R})$. By the Laplace formula [HvNVW18, Proposition G.4.1 p. 532] for the resolvent, we have for $\operatorname{Re}\lambda>0$ the equality
@@
and can be seen by [HvNVW16, Proposition 5.2.2 p. 389] as a Fourier multiplier with symbol $\mathrm{i}\,\mathrm{sign}(\xi)$, i.e., we have
(8.51) $\widehat{\mathcal{H}(f)}(\xi)=-\mathrm{i}\,\mathrm{sign}(\xi)\hat{f}(\xi),\quad f\in\mathcal{S}(\mathbb{R}),\xi\in\mathbb{R}.$
@@
In dimension $d=2$, we have $N=2^{\lfloor\frac{d}{2}\rfloor}=2$ and we can consider the Pauli matrices $\sigma_{1}\overset{\mathrm{def}}{=}\begin{bmatrix}0&1\\ 1&0\end{bmatrix}$ and $\sigma_{2}\overset{\mathrm{def}}{=}\begin{bmatrix}0&-\mathrm{i}\\ \mathrm{i}&0\end{bmatrix}$. Then we can write the Euclidean Dirac operator on $\mathbb{R}^{2}$ as
[…]
where we use the classical operators
$\partial_{\overline{z}}\overset{\mathrm{def}}{=}\frac{1}{2}\big(\partial_{x}+\mathrm{i}\partial_{y}\big)\quad\text{and}\quad\partial_{z}\overset{\mathrm{def}}{=}\frac{1}{2}(\partial_{x}-\mathrm{i}\partial_{y}).$
See [LaM89, p. 119] for more information on this operator. In the sequel, we consider the domain $\operatorname{dom}D=\mathrm{W}^{1,p}(\mathbb{C})\oplus\mathrm{W}^{1,p}(\mathbb{C})$ of the Banach space $\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})$. The square of this operator is given by
[…]
Following (8.25), we can introduce the isometric homomorphism $\pi\colon\mathrm{L}^{\infty}(\mathbb{C})\to\mathrm{B}(\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C}))$, $f\mapsto\begin{bmatrix}M_{f}&0\\ 0&M_{f}\\ \end{bmatrix}$.
Remark 8.30
Recall that $\Delta=4\partial_{\overline{z}}\partial_{z}$ The operators $\partial_{\overline{z}}$, $\partial_{z}$ and $\Delta$ can be seen as unbounded Fourier multipliers on $\mathrm{L}^{p}(\mathbb{C})$ with symbol $\mathrm{i}\pi\zeta$, $\mathrm{i}\pi\overline{\zeta}$ and $|\zeta|^{2}$, see [AIM09, p. 99].
Proposition 8.13 gives the following result.
@@
Suppose that $1<p<\infty$. Then
@@
(8.61) $D,f]=\frac{2}{\mathrm{i}}\begin{bmatrix}0&-M_{\partial_{\overline{z}}f}\\ M_{\partial_{z}f}&0\end{bmatrix}$
@@
Proof : We have
@@
The inequality $\big\|[D,f]\big\|\leqslant 2\max\Big\{\left\|\partial_{z}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})},\left\|\partial_{\overline{z}}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})}\Big\}$ follows from a simple computation. Indeed, for any $u,v\in\mathrm{L}^{p}(\mathbb{C})$ we have
@@
From Theorem 8.18, we obtain the following result.
Theorem 8.32
@@
Note that we have